### v2

# EVQAScore: Efficient Video Question Answering Data Evaluation

Abstract
Video question-answering (QA) is a core task in video understanding. Evaluating the quality of video QA and video caption data quality for training video large language models (VideoLLMs) is an essential challenge. Although various methods have been proposed for assessing video caption quality, there remains a lack of dedicated evaluation methods for Video QA. To address this gap, we introduce EVQAScore, a reference-free method that leverages keyword extraction to assess both video caption and video QA data quality. Additionally, we incorporate frame sampling and rescaling techniques to enhance the efficiency and robustness of our evaluation, this enables our score to evaluate the quality of extremely long videos. Our approach achieves state-of-the-art (SOTA) performance (32.8 for Kendall correlation and 42.3 for Spearman correlation, 4.7 and 5.9 higher than the previous method PAC-S++) on the VATEX-EVAL benchmark for video caption evaluation. Furthermore, by using EVQAScore for data selection, we achieved SOTA results with only 12.5% of the original data volume, outperforming the previous SOTA method PAC-S and 100% of data. Our code is available at https://anonymous.4open.science/r/EVQAScore-FFA6/.
1 Introduction
With the rapid advancements in large language models (LLMs) [30, 36] and multimodal large language models (MLLMs) [45, 40],
Figure 1: We give an example of EVQAScore, from the figure we can see that Answer1 is low quality and contains few information while Answer2 contains more information and is a more comprehensive answer. We can see our EVQAScore is the only score can correctly identify the high quality data.
Among MLLMs, VideoLLMs achieve competitive performance in traditional multimodal tasks such as visual recognition [44], video understanding [41, 35], and action recognition [39]. Moreover, their excellent language understanding capabilities enable strong performance in text-rich tasks, such as video question-answering [9] and video-centric dialogues [39].
Most existing VideoLLMs focus on modifying model architectures to effectively utilize information from multiple modalities [38, 39, 26, 43, 17]. While model effectiveness is crucial, data quality also plays a significant role in the success of VideoLLMs. For instance, Li et al. [18], Wang et al. [39] demonstrate that higher-quality training data enhances the performance of VideoLLMs. Therefore, evaluating video QA data quality is essential. However, previous methods such as EMScore and PAC-S [34, 33] focus exclusively on video caption evaluation while neglecting video QA evaluation. Moreover, these methods struggle with understanding the correct semantics of longer captions or QA pairs. As illustrated in Figure 1, EMScore and PAC-S fail to identify higher-quality video QA pairs. These methods overlook the importance of efficient and robust data management and face three key challenges:
C1. Inability to Accurately Understand Caption and QA Semantics Previous methods such as EMScore [34] and PAC-S [33] rely on TF-IDF to evaluate video caption data. However, TF-IDF merely computes word frequency and does not capture semantic meaning. As a result, these methods often fail to accurately evaluate video captions and are unable to assess video QA data with precision. Thus, there is a need for improved semantic understanding in Video QA evaluation.
C2. Low Evaluation Efficiency Given the extensive volume and length of video data, large-scale training datasets frequently occupy substantial storage space, ranging from several hundred gigabytes to tens of terabytes [38, 39, 20]. Evaluating video caption and QA data is time-consuming and computationally costly. Previous methods like EMScore [34] and PAC-S [33] process every frame, leading to considerable computational waste. Therefore, developing a computationally efficient evaluation metric is essential.
C3. Lack of VideoLLM Experiments Data plays a crucial role in the success of VideoLLMs [38, 39, 26, 43, 17]. Although various methods have been proposed to evaluate the quality of video caption data, they have not incorporated VideoLLM experiments to assess the effectiveness of their methods in training VideoLLMs. Consequently, additional VideoLLM experiments are needed.
To address these challenges, we introduce EVQAScore, which utilizes LLMs to extract keywords for enhanced video captioning and video QA data evaluation. To support the evaluation of longer videos, we efficiently compute EVQAScore through uniform frame sampling, which reduces computational costs by a factor of 30 without affecting results. Compared to previous methods such as EMScore [34] and PAC-S [33], our approach can be applied to video QA evaluation and achieves superior visual understanding through keyword extraction, outperforming the TF-IDF technique used by EMScore and PAC-S. Additionally, due to its efficiency and keyword extraction capabilities, our method can handle extremely long captions and videos, a feat not possible for EMScore and PAC-S because of the 77-word limit imposed by CLIP. The core contributions of this paper are summarized as follows:
  • New Method: We introduce a novel approach, EVQAScore, designed for the evaluation of Video QA and video caption data quality. To the best of our knowledge, this is the first systematic effort to evaluate the quality of Video QA data. Our approach leverages LLMs and keyword extraction techniques to improve the understanding of QA data in Video QA evaluation. Additionally, we employ uniform frame sampling to enhance efficiency and scalability, allowing EVQAScore to be applied to extremely long videos and large datasets.
  • Efficient Video QA Data Evaluation: We use uniform frame extraction, resulting in a computational cost reduction of over 30 times compared to processing all frames. Furthermore, we demonstrate that this reduction does not lead to any performance degradation.
  • SOTA Performance for Video Caption Evaluation: We utilize LLM for keyword extraction for better caption and QA understanding. We evaluated the performance of our method on the VATEX-Eval benchmark for Video Caption data, achieving a Kendall correlation of 32.8 and a Spearman correlation of 42.3, which are 4.7 and 5.9 points higher than the previous method PAC-S++.
  • SOTA Performance for Video QA Evaluation: We constructed a dataset comprising 400K video entries, consisting of 200K noisy samples and 200K high-quality samples. Using our EVQAScore and the previous SOTA model PAC-S, we filtered the data, retrieving a significantly lower amount of noisy data than PAC-S. As shown in Figure 1, our method is the only method that can identify low-quality video QA data. Additionally, when evaluated on ActivityNet, MSRVTT, MSVD, TGIF, MVBench, and two longer benchmarks—VideoChatGPT Bench and VideoChatGPT Bench Diverse—our EVQAScore outperformed PAC-S and 100% of data with only 12.5% of data.
2 Related Work
2.1 Video Data Evaluation
Traditional image and video captioning evaluations have relied on metrics like BLEU[32], METEOR[4], ROUGE[21], CIDEr[37], and SPICE[1]. These metrics, based on n-gram matching, primarily focus on lexical similarity between generated and reference captions. While effective in text-based tasks like machine translation, they struggle to capture the semantic richness of video content, leading to poor correlation with human judgments, especially when evaluating dynamic, time-sensitive visual elements in videos.
Recently, methods like CLIP-Score[8] leverage pre-trained vision-language models to measure the coarse-grained alignment between video frames and captions. However, CLIP-Score lacks fine-grained analysis. To address this, Emscore[34] introduced fine-grained matching between video frames and specific words in the captions, significantly improving semantic alignment. Building upon Emscore, PAC-Score[33] further enhances the precision of video-text alignment by incorporating a positive-augmented CLIP model, leading to a more comprehensive evaluation of video content.
2.2 Data Quality and Selection
The advent of large language models has brought about a substantial increase in the volume of training data [36, 31]. In this scenario, the quality and quantity of data become paramount [3]. In this section, we mainly discuss the effectiveness of data quality and data selection.
Data Quality
: High-quality data can significantly enhance the performance of models [29]. As the volume of data increases, ensuring high data quality becomes more challenging because it requires more resources for data cleaning, selection and annotation [3]. Poor quality data can lead to models learning incorrect patterns and making inaccurate predictions.
Data Selection
: LLMs-based methods were commonly used in data selection [3]. For instance, Du et al. [6] leverages DeBERTa [7] for scoring, retaining high-quality data, and combining it with the k-center greedy algorithm to select diverse data. Chen et al. [5] score the accuracy of data using ChatGPT to pick out high-quality data. Xu et al. [42] use GPT-4 to rewrite data to increase their complexity and then streamline it by reducing its variety and improving its quality.
2.3 Video Large Language Models.
Recently, inspired by the remarkable understanding capabilities of LLMs [14], multimodality breakthroughs [14, 15, 11, 12, 16, 13, 25, 22, 10], and pre-trained models, researchers have started using LLMs to understand videos, achieving SoTA results [18, 43, 2, 39, 20]. VideoLLaMA [43] is one of the pioneering studies in VideoLLMs, utilizing a visual encoder and a video Q-Former projector to understand videos. However, due to its Q-Former structure, the computational cost is high. To address this, subsequent works [20] adopted the LLaVA [24, 23] MLP structure, significantly reducing computational costs while still achieving SoTA performance. Similarly, MiniGPT4Video [2] uses an MLP adapter for efficient training.
Another notable series of models includes VideoChat, VideoChat2, InternVideo, and InternVideo2 [17, 18, 38, 39]. These models utilize an enormous amount of data to train a transformer-structured adapter, achieving SoTA performance. By leveraging large-scale datasets and advanced transformer architectures, these models excel in comprehending and processing multimodal video content, further pushing the boundaries of video understanding capabilities.
3 Method
Figure 2: Illustration of the computation of EMScore which utilize coarse to fine video evaluation. We utilize keyword extraction and then use frames and keywords as ground truth for precision and recall calculation.
3.1 Motivation and problem formulation
Inspired by EMScore and PAC-S, we extend these video-caption data quality evaluation methods to assess the quality of VideoQA data. We aim to address three key limitations of existing methods.
Difficulty in Evaluating Long Videos:
Previous methods, such as PAC-S, face challenges when dealing with longer videos due to the high computational costs and unable to understand long captions and QAs due to the CLIP token limitation (77 tokens). To address this issue, we employ uniform frame extraction to reduce computational costs and utilize keyword extraction for better QA and caption understanding.
Inability to Capture Accurate Semantic Meaning of Captions:
Methods like PAC-Score and EMScore rely on TF-IDF for semantic understanding, which is limited because it only computes word frequency and cannot capture contextual semantics. Therefore, we utilize LLMs to extract keywords for more accurate understanding of QA and caption semantics.
Evaluating VideoQA Data for VideoLLMs:
Video QA is increasingly important in areas such as SFT, requiring adaptable methods capable of handling these tasks. As such, we apply EVQAScore for Video QA data filtering.
Problem Formulation
Given a video $V$ and a corresponding question-answer pair $(Q,A)$, our goal is to evaluate the quality of the VideoQA data using our proposed evaluation metric, EVQAScore. This metric assesses the data’s suitability for training VideoLLMs by scoring based on several criteria, including semantic relevance and informativeness of the answer $A$ with respect to the video $V$ and the question $Q$.
Formally, for a given data instance $(V,Q,A)$, we compute EVQAScore to evaluate the quality of $(V,Q,A)$:
$\text{EVQAScore}(V,Q,A)$
Our objective is to filter and select high-quality data based on the EVQAScore to improve the performance and training efficiency of VideoLLMs.
3.2 Efficiency of EVQAScore
Keyframe Extraction
To address the challenge of evaluating long videos efficiently, we adopt a uniform frame sampling strategy to reduce the number of frames processed while retaining critical information. By sampling keyframes at regular intervals, we capture essential content and minimize computational costs. The frame sampling process is defined as:
$S=\{v_{i}\mid i=k\cdot l,\;k\in\mathbb{Z}_{\geq 0},\;k\cdot l\leq m\}$ (1)
where $S$ represents the set of sampled frames, $v_{i}$ denotes the $i$-th frame in the video $V$, $m$ is the total number of frames, and $l$ is the sampling interval. This approach reduces redundancy while ensuring a diverse representation of keyframes.
In our experiment we select sampling interval $l$ to be 10, 20, and 30. We present the result in Table 1.
Keywords Extraction
Given a text input $X$, we use a large language model (LLM) to extract a list of key phrases $\{k_{1},k_{2},\ldots,k_{n}\}$ as follows:
$\{k_{1},k_{2},\ldots,k_{n}\}=\text{LLM}_{\text{extract}}(X,P)$
$K=\{k_{1}\oplus k_{2}\ldots\oplus k_{n}\}$
where $\text{LLM}_{\text{extract}}$ denotes the language model’s extraction function, and $P$ represents the prompt designed to guide the extraction of relevant phrases from the input $X$. This approach enhances the efficiency of video-text alignment by focusing on the most relevant key phrases identified through the prompt-based extraction.
3.3 Video caption evaluation
The improved scoring method now combines the coarse-grained score and the modified fine-grained score, which includes keyframe-based matching and keyword-based matching. This enhanced PAC-Score improves efficiency when handling long videos and complex text descriptions, focusing on the most essential frames and phrases.
Coarse-Grained EVQAScore
Coarse-grained score of $\text{EVQAScore}(V,X)_{c}$ is:
$\text{EVQAScore}(V,X)_{c}=f_{V}^{\top}f_{X}$
where $f_{V}$ is the embedding of the video, which is calculated by pooling the feature of video frames in sampled frames $S$, respectively. $f_{X}$ is the feature of caption or QA pair.
Fine-Grained EVQAScore
First, the precision ($P$) and recall ($R$) for the caption $X$ are updated using their keywords as follows:
$P(V,X)_{f}=P\left(S,K\right)_{f}=\frac{1}{|X|}\sum_{k_{j}\in K}\max_{v_{i}\in S}f_{v_{i}}^{\top}f_{k_{j}}$
$R(V,X)_{f}=R\left(S,K\right)_{f}=\frac{1}{|V|}\sum_{v_{i}\in S}\max_{k_{j}\in K}f_{v_{i}}^{\top}f_{k_{j}}$
where $f_{v_{i}}$ and $f_{k_{j}}$ represent the CLIP embedding scores of frame $v_{i}$ and keyword $k_{j}$. Subsequently, the fine-grained EVQAScore is calculated as:
$\text{EVQAScore}(V,X)_{f}=2\frac{P(V,X)_{f}\cdot R(V,X)_{f}}{P(V,X)_{f}+R(V,X)_{f}}$
Finally, the overall EVQAScore is computed by combining the original coarse-grained score with the updated fine-grained score, which incorporates both keyframe and keyword matching.
$\scriptsize\text{EVQAScore}(V,X)=\frac{\text{EVQAScore}(V,X)_{c}+\text{EVQAScore}(V,X)_{f}}{2}$ (2)
This formula ensures that both visual and textual representations are aligned more effectively, improving precision, recall, and computational efficiency by focusing on keyframes and keywords.
3.4 Video QA Evaluation
We compute EVQAScore for Video QA data using the following formula:
$\text{EVQAScore}(\text{Concat}(Q,A),V)\cdot\log(n+1)$
In this formulation, $\text{Concat}(Q,A)$ represents the concatenation of the question $Q$ and the answer $A$, while $n$ denotes the number of extracted keywords. We posit that a higher number of extracted keywords often indicates a richer and more informative caption, thereby warranting a higher score. We show the advancement of our EVQAScore in Section 4.
Table 1: Comparison of time spent managing video data between uniform keyframe extraction (every 10 frames) and no keyframe extraction for 1000 videos. Extracting keyframes leads to a significantly shorter time.
Method Time (s) Kendall Spearman
Uniform (Interval=30) 464 25.70 33.41
Uniform (Interval=20) 927 25.72 33.43
Uniform (Interval=10) 1857 25.73 33.44
None 18504 25.67 33.37
Table 2: Performance Comparison on VATEX-EVAL Dataset (No Reference). The table presents the Kendall $\tau_{b}$ and Spearman $\rho$ values, comparing Emscore, PAC-Score, and our proposed Emscore+. Emscore+ achieves the best performance, as indicated by the improvement values in parentheses.
Metric Kendall $\tau_{b}$ Spearman $\rho$
Emscore 23.2 30.3
PAC-S 25.1 32.6
EVQAScore(PAC-S) 25.6 33.3
(+0.5) (+0.7)
PAC-S++ 28.1 36.4
EVQAScore(PAC-S++) 32.8 42.3
(+4.7) (+5.9)
Table 3: Performance of VideoLLaVA and VideoLLaMA2 on ActivityNet, MSRVTT, MSVD, and TGIF shows that with only 12.5% of the data, EVQAScore outperforms the baseline using 100% of the data. Additionally, our approach surpasses PAC-S on both models across all evaluation datasets.
Method Model Data Proportion ActivityNet MSVD MSRVTT TGIF
Accuracy Score Accuracy Score Accuracy Score Accuracy Score
Baseline VideoLLaMA2 100% 28.56 1.36 40.04 1.99 36.37 1.80 32.13 1.59
PAC-S 25% 41.17 2.07 71.46 3.53 62.17 3.08 57.92 2.87
PAC-S 12.5% 45.41 2.28 81.31 4.04 62.91 3.12 57.00 2.82
EVQAScore 25% 53.41 2.62 82.03 4.07 69.76 3.45 57.39 2.83
EVQAScore 12.5% 57.05 2.82 74.67 3.72 69.03 3.43 58.58 2.90
Baseline VideoLLaVA 100% 15.64 0.78 19.93 1.00 24.97 1.24 11.72 0.58
PAC-S 25% 32.27 1.60 46.11 2.30 50.13 2.49 41.36 2.05
PAC-S 12.5% 35.41 1.77 52.13 2.60 53.23 2.65 43.33 2.14
EVQAScore 25% 41.56 2.05 58.84 2.93 57.98 2.88 43.87 2.15
EVQAScore 12.5% 44.42 2.33 57.32 2.86 56.16 2.79 37.18 1.83
4 Experiments
In this section, we employ our Keyprompt method on both short and long benchmarks. We aim to address the following questions to verify the effectiveness of our method:
Q1: Can our EVQAScore achieve efficient computation without compromising evaluation performance?
Q2: Does our EVQAScore outperform all baseline methods in video caption data evaluation?
Q3: Does our EVQAScore outperform all baseline methods in selecting Video QA pairs?
Q4: Does our EVQAScore demonstrate superior performance compared to baseline methods when training VideoLLMs?
Q5: Can we visualize the advantages of our EVQAScore relative to previous video caption evaluation methods?
4.1 Experimental Setting
Datasets.
For the QA evaluation, we selected 200K samples from two datasets: the VideoInstruct100K [27] dataset and the VCG+ [28] 112K dataset. We then augmented these 200K samples with an additional 200K low-quality (noisy) data entries created by substituting each answer with a single keyword extracted from the answer sentence.
Evaluation Metrics
To assess the alignment of our metric with human judgment, we utilized the VATEX-EVAL dataset to evaluate the correlation between our scores and human ratings in video captioning tasks. The VATEX-EVAL dataset is specifically designed to measure the correlation between automatic evaluation metrics and human judgment. It consists of 3,000 validation videos from the VATEX dataset, with a total of 18,000 candidate captions, six for each video.
For Video QA ability evaluation, following VideoChatGPT [26, 28], we evaluated the ActivityNet, MSRVTT, MSVD, and TGIF datasets using LLaMA-3.1-Instruct to compare the model’s answers against the ground-truth captions and provide a score from 0 to 5 for each caption. Additionally, we used VideoChatGPT Bench and VideoChatGPT Bench Diverse [26, 28] to evaluate model performance on longer videos. Furthermore, we employed MVBench [19], a comprehensive benchmark, to evaluate the performance of our method on VideoLLMs.
Models.
We employed the positive-augmented CLIP visual encoder [33] to extract embeddings from individual frames, applying average pooling to represent the entire video. Similarly, we used the corresponding textual encoder to generate token embeddings, producing a comprehensive caption representation. Additionally, we utilized the LLaMA3.1-8B-Instruct model to identify key terms from question-answer pairs, highlighting the essential elements of the dialogue.
For VideoLLMs, we selected two previous SOTA model Video-LLaVA [20] and Video-LLaMA2 [43] to evaluate the quality of our selected data.
Baselines.
We used EMScore, PAC-S, and PAC-S++ as baselines. While all three employ the same computational formula, they differ in the specific CLIP models utilized for their calculations.
Settings.
For LLaMA3.1-8B-Instruct, we applied the official hyperparameters from the repository. For the CLIP model, we used the clip-ViT-B-32 variant for keyframe selection. All experiments were conducted on a machine equipped with 8 NVIDIA H800 GPUs, a 192-core CPU, and 960GB of memory.
Table 4: Comparison of Noisy data filtered by PAC-S and our EVQAScore. We select much less noisy data.
Extract 25% data
Method noisy videochatgpt videochatgpt+ total
PAC-S 27928 40012 33764 101704
EVQAScore 1101 53341 47262 101704
Extract 12.5% data
PAC-S 10832 21095 18925 50852
EVQAScore 450 25642 25165 50852
4.2 Computational Efficient EVQAScore
To address Q1, we conduct uniform frame extraction in EVQAScore. We select the sampling interval to be 10, 20 and 30 frames. The results is presented in Table 1, they illustrate that using uniform keyframe extraction substantially reduces the time required for managing video data, with significantly shorter processing times compared to not extracting keyframes. For instance, applying a uniform interval of 30 frames results in a processing time of 464 seconds, compared to 18,504 seconds for no keyframe extraction, representing a drastic time reduction. Importantly, this improvement in efficiency comes with nearly no impact on performance, as evidenced by the Kendall and Spearman correlation scores, which remain virtually unchanged across different intervals. This demonstrates that uniform frame selection is a highly efficient approach, achieving considerable time savings without compromising the evaluation quality.
4.3 EVQAScore for Video Caption Evaluation
To address Q2, we conducted video caption evaluation experiments on the VATEX-EVAL dataset. The results are presented in Table 2. Our findings demonstrate that EVQAScore consistently outperforms both EMScore and PAC-S in correlating with human judgment. Specifically, EVQAScore achieves a Kendall $\tau_{b}$ score of 25.6 and a Spearman $\rho$ score of 33.3 in the No Reference setting using the PAC-S CLIP model, representing an improvement of 0.5 points in Kendall and 0.7 points in Spearman over PAC-Score. Additionally, further experiments using PAC-S++ checkpoints resulted in SOTA performance, with a Kendall $\tau_{b}$ score of 32.8 and a Spearman $\rho$ score of 42.3, marking improvements of 4.7 points and 5.9 points over the PAC-S++ baseline. These results indicate that our proposed method better aligns with human perception of video-caption consistency in the No Reference evaluation setting.
This strong performance highlights EVQAScore’s capacity to enhance existing scoring systems. It demonstrates that EVQAScore excels not only in evaluating Video QA data but also in video caption data evaluation. This further underscores the potential of EVQAScore for high-fidelity video caption data assessment. As our method utilizes only keywords, previous caption-based methods may suffer from the token limitation (77 tokens) of CLIP, making our approach more suitable for evaluating very long videos and their comprehensive caption. This strong performance highlights EVQAScore’s capacity to amplify existing scoring systems, further emphasizing its potential for high-fidelity video QA data evaluation.
4.4 Video QA Data Filtering
Table 5: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench-Diverse.
Method Model Data Proportion Video Understanding Metrics High-Level Metrics
CI DO CU TU CO Avg. Caption Spatial Reasoning
Baseline VideoLLaMA2 100% 0.31 0.15 0.49 0.14 0.37 0.29 0.01 0.44 0.24
PAC-S 25% 25% 1.56 1.87 2.17 0.93 1.82 1.67 0.79 1.66 2.70
PAC-S 12.5% 12.5% 1.57 1.80 2.10 1.01 1.87 1.67 0.43 1.76 2.87
EVQAScore 25% 25% 1.83 2.39 2.57 1.09 2.33 2.04 1.32 1.92 3.09
EVQAScore 12.5% 12.5% 1.89 2.51 2.64 1.07 2.33 2.08 1.33 2.04 3.24
Baseline VideoLLaVA 100% 0.33 0.16 0.52 0.13 0.36 0.3 0.01 0.45 0.28
PAC-S 25% 25% 1.63 1.92 2.20 0.87 1.79 1.68 0.75 1.78 2.81
PAC-S 12.5% 12.5% 1.62 1.86 2.15 0.93 1.80 1.67 0.56 1.79 2.87
EVQAScore 25% 25% 1.75 2.27 2.43 0.86 2.07 1.88 1.05 1.94 3.04
EVQAScore 12.5% 12.5% 1.76 2.39 2.47 0.83 1.97 1.88 1.00 1.98 3.25
Table 6: Performance of VideoLLaVA and VideoLLaMA2 on VCGBench.
Method Models Data Proportion CI DO CU TU CO Avg.
Baseline VideoLLaMA2 100% 0.60 0.29 0.88 0.21 1.05 0.61
PAC-S 25% 2.02 1.70 2.35 1.59 2.48 2.03
PAC-S 12.5% 2.39 2.15 2.75 1.73 2.75 2.35
EVQAScore 25% 3.01 2.98 3.38 2.33 3.05 2.95
EVQAScore 12.5% 3.02 3.14 3.39 2.23 3.09 2.97
Baseline VideoLLaVA 100% 0.59 0.30 0.87 0.20 0.96 0.58
PAC-S 25% 2.18 1.96 2.53 1.60 2.54 2.16
PAC-S 12.5% 2.24 2.08 2.63 1.82 2.60 2.27
EVQAScore 25% 2.77 2.82 3.15 2.16 3.04 2.79
EVQAScore 12.5% 2.68 2.93 3.09 2.15 2.71 2.71
Table 7: Performance of VideoLLaVA and VideoLLaMA2 on MVBench using filtered data by PAC-S and EVQAScore.
Method Model Data Proportion Avg.
Baseline VideoLLaMA2 100% 34.09
PAC-S 25% 25% 31.43
PAC-S 12.5% 12.5% 34.24
EVQAScore 25% 25% 34.89
EVQAScore 12.5% 12.5% 37.60
Baseline VideoLLaVA 100% 30.68
PAC-S 25% 25% 37.10
PAC-S 12.5% 12.5% 37.53
EVQAScore 25% 25% 38.01
EVQAScore 12.5% 12.5% 38.96
To address Q3, we perform data filtering using our EVQAScore metric alongside the previous state-of-the-art (SOTA) data evaluation baseline, PAC-S. The results presented in Table 4 illustrate the effectiveness of the EVQAScore metric in filtering Video QA data. For both the top 25% and 12.5% data extractions, EVQAScore significantly reduces noisy data while retaining a higher proportion of VideoChatGPT and VideoChatGPT+ data compared to the PAC-S method. Specifically, EVQAScore retains only 1,101 and 450 noisy data points for the 25% and 12.5% extractions, respectively, while preserving a greater volume of high-quality data. This indicates that EVQAScore is more selective, thereby enhancing data quality and processing efficiency for Video QA tasks.
By reducing the amount of low-quality data, we are able to train more effective VideoLLMs. Further experiments presented in Section 4.5 demonstrate the impact of EVQAScore on VideoLLM performance.
4.5 Performance of Video-QA Evaluation
To address Q4, we train two models VideoLLaVA and VideoLLaMA2 on our filtered datasets. Then we conduct comprehensive evaluation on several benchmarks, i.e. Activitynet, MSVD, MSRVTT, TGIF, MVBench, VCGBench and VCGBench Diverse. Our EVQAScore achieves SOTA performance across diverse models and benchmarks.
SOTA Performance Across Different Benchmarks
The summarized results for ActivityNet, MSVD, MSRVTT, and TGIF are presented in Table 3. It is evident that our EVQAScore achieves state-of-the-art (SOTA) performance on all benchmarks for both VideoLLaVA and VideoLLaMA2 models. Notably, EVQAScore demonstrates the ability to enhance long-video performance, as seen with ActivityNet (average video duration of 113.74s), as well as short-video performance, exemplified by TGIF (average duration of 11.93s). Furthermore, we extended the evaluation by utilizing our trained model on three more comprehensive benchmarks. The results for VCGBench are summarized in Table 6, for VCGBench-Diverse in Table 5, and for MVBench in Table 7. Across all these datasets, our method consistently outperforms previous approach PAC-S. Moreover, our EVQAScore outperforms the 100% data baseline in all benchmarks, this shows the importance of evaluating data quality and low quality data filtering.
SOTA Performance Across Different Models
Figure 3: A case study on evaluating very long QA pairs demonstrates that EVQAScore can effectively assess such cases, whereas CLIPScore, EMScore, and PAC-S struggle to do so.
We selected two different models: VideoLLaVA and VideoLLaMA2. These models have distinct architectures and use different encoders and foundation models. Specifically, VideoLLaVA employs LanguageBind as its encoder and Vicuna 1.5 as the foundation model, while VideoLLaMA2 utilizes CLIP as its encoder and LLaMA2 as the foundation model. Across the four benchmarks—ActivityNet, MSVD, MSRVTT, and TGIF presented in Table 3, VCGBench in Table 6, VCGBench-Diverse in Table 5, and MVBench in Table 7—our filtered data outperformed the previous SOTA baseline, PAC-S, for both model architectures. This demonstrates the robustness of our EVQAScore.
In this section, we demonstrate that our method effectively filters high-quality data to enhance VideoLLM performance. We showed that EVQAScore can improve the performance of models with different architectures across all benchmarks, underscoring the wide application potential of our EVQAScore.
4.6 Case Study: Evaluating Video Captions
To address Q5, we present case study, we demonstrate the effectiveness of EVQAScore using two cases.
EVQAScore can correctly evaluate comprehensive answers
We summarize the first case in Figure 1. Given the question: ”What is the man doing in the video?” the first answer contains only one word and is of low quality, while the second answer is longer and of higher quality. Compared with CLIPScore (using the pooling feature of video), EMScore, PAC-S, and EVQAScore, the video QA pair indicates that our method is the only approach capable of recognizing the second answer as being of higher quality. This demonstrates the robustness of our method in evaluating Video QA data.
EVQAScore can evaluate longer QAs
We summarize the second case in Figure 3. From the figure, we observe a man playing bagpipes. When asked to describe the content of the video, the first answer, ”pipes,” is too short and incorrect. In contrast, the second answer is comprehensive and considerably longer. However, due to the absence of keyword extraction, methods such as CLIPScore (using the pooling feature of video), EMScore, and PAC-S suffer from the limited token count (77 tokens) imposed by CLIP, requiring them to truncate the caption, thereby preventing a full understanding of the QA. Consequently, CLIPScore, EMScore, and PAC-S fail to demonstrate that Answer 2 outperforms Answer 1.
In this section, we use two real cases to illustrate the effectiveness of EVQAScore. In recent years, understanding longer videos has become increasingly essential, and evaluating the quality of long videos is a critical task. When applied to longer videos—often spanning hours and involving thousands of words in QA or captions—our method demonstrates a higher potential for success.
5 Conclusion
In this work, we address the critical challenge of evaluating video QA data quality by introducing EVQAScore, a reference-free evaluation method that leverages keyword extraction to assess both video caption and video QA data quality. By incorporating frame sampling and rescaling techniques, our method effectively handles the evaluation of extremely long videos, enhancing efficiency and robustness. EVQAScore achieves state-of-the-art performance on the VATEX-EVAL benchmark, with scores of 32.8 for Kendall correlation and 42.3 for Spearman correlation in video caption evaluation. Moreover, by applying EVQAScore for data selection, we demonstrate its effectiveness in selecting much less noisy data. Moreover, we achieved SOTA results, surpassing the previous SOTA method PAC-S and 100% data with only 12.5% of the original data volume, highlighting the potential for substantial improvements in data efficiency and overall video understanding.
Supplementary Material
6 Keywords Extraction Prompt
Figure 4: We provide the prompt used for keywords extraction